OpenStudy (anonymous):

Simplify. x / 6x-x^2 Please help me!!

OpenStudy (anonymous):

-1/x-6

OpenStudy (anonymous):

That's not an option. The options are: 1/6-x where x=0,6 1/6-x where x=6 1/6x where x = 0 1/6

OpenStudy (anonymous):

is it x/(6x-x^2)

OpenStudy (anonymous):

or is the -x^2 not part of the fraction

OpenStudy (anonymous):

1/6-x where x=6

OpenStudy (anonymous):

It is x over 6x-x^2 this is the entire fraction

OpenStudy (irishboy123):

lmLtfy \( \frac{x}{ 6x-x^2} \) so try losing an 'x', ie divide top and bottom by x

OpenStudy (anonymous):

So, I would get 1/6-x ; where x= 6 ????

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

correct

OpenStudy (irishboy123):

\(\frac{x}{6x-x^2} = \frac{x}{x}.\frac{1}{6-x} = \frac{1}{6-x}\) well done

OpenStudy (anonymous):

you're sure, that I am right? answer says: \[1 / 6-x ; x \neq 6\]

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (irishboy123):

yes

OpenStudy (anonymous):

I posted another question! I really need help. I do not understand this